OpenStudy (anonymous):

the graph of the following system of equations is -2x+y=3 4x+2y=2 a.)overlapping lines b.)parallel lines c.)intersecting lines

OpenStudy (studygurl14):

Convert to y = mx + b form -2x + y = 3 y = 2x + 3 4x + 2y = 2 2y = -4x + 2 y = -2x + 1 They can't be parallel, right? Why? Because they have different slopes. They can't be overlapping, because they aren't the same exact equations. Thus, they must be intersecting

OpenStudy (anonymous):

can you help me with another one?

OpenStudy (studygurl14):

Sure